# Random Time-hopping Secure Ranging Strategy Against Distance-Reduction Attacks in UWB

### v1

@@ III-C Performance Analysis
Generally it can be assumed that $t$ follows the uniform distribution $\text{U}(\text{T}_{\text{min}},\text{T}_{\text{max}})$. Taking into consideration (5) and (6), the attack success probability ${\mathbb{P}_{\text{s}}}^{{}^{\prime}}(t)$ at the time offset $t$ becomes:
${\mathbb{P}_{\text{s}}}^{{}^{\prime}}(t)$ $={\mathbb{P}_{\text{s}}}(t)~{}\mathbb{P}({T_{\text{SFD}}<t<T_{\text{payload}}})$ (7)
$=2F_{\text{BIN}}({0.5\text{N}-0.5\theta{(|{p(t)}|}^{2})^{-1}|{\text{N},0.5}})\frac{\Delta t_{0}}{\Delta t_{1}},$
@@
$T_{\text{SFD}}<t-\Delta t<T_{\text{payload}}.$ (8)
Let the random variable $Y=t-\Delta t$, the attack success probability ${\mathbb{P}_{\text{s}}}^{{}^{\prime\prime}}(t)$ after adopting the random time-hopping mechanism becomes:
${\mathbb{P}_{\text{s}}}^{{}^{\prime\prime}}(t)$ $={\mathbb{P}_{\text{s}}}(t)~{}\mathbb{P}({T_{\text{SFD}}<t-\Delta t<T_{\text{payload}}})$ (9)
$=\left(2F_{\text{BIN}}({0.5\text{N}-0.5\theta{(|{p(t)}|}^{2})^{-1}|{\text{N},0.5}})\right.$
×$\left.\times{\int_{T_{\text{SFD}}}^{T_{\text{payload}}}f_{Y}}(y)dy\right),$
where ${f_{Y}}(y)$ is the probability density function (pdf) of the random variable $Y$, which is formed by the convolution of two uniform distributions.
Let $\Delta t_{2}=T_{\text{max}}^{\text{hop}}-T_{\text{min}}^{\text{hop}}$, then ${f_{Y}}(y)$ is an isosceles trapezoid between $T_{\text{max}}^{\text{hop}}$ and $T_{\text{min}}^{\text{hop}}$, with its height and upper base $\frac{1}{\Delta t_{2}}$ and $\Delta t_{2}-\Delta t_{1}$ respectively. Since the range of delay $\Delta t_{2}$ for the random time-hopping proposed in this paper is much larger than the original time interval $\Delta t_{1}$ (i.e. $\Delta t_{2}\gg\Delta t_{1}$), ${f_{Y}}(y)$ can be approximated as a rectangle with bottom $\Delta t_{2}$ and height $\frac{1}{\Delta t_{2}}$. Therefore, the reduction factor $G$ of the attack success probability (the gain of this strategy) is:
@@ IV-A Numerical Simulation Evaluation
The signal parameter configurations for simulation are detailed in Table I.
@@
Payload encoding RS &convolution RS &convolution
@@
The processing flow of the receiver used in this paper includes shaping filtering, downsampling, cross-correlation, timestamp estimation and data demodulation utilizing a RAKE receiver architecture. The length of BTW is fixed at 400 samples (at a sampling rate of 2 GHz). MPEP and PAPR are set to 0.5 and 2, respectively. Moreover, the distance between two legitimate devices is 10 m and an attack is considered successful when the measurement result is less than 5 m. 20,000 DS-TWR simulations under the Ghost Peak attack are performed for each case.
SIR is used to denote the ratio between the power of the STS segment of the legitimate signal and that of the attack signal, and Signal-to-Noise Ratio (SNR) in the environment is set to -10 dB. The attack success probability simulated under different SIR and attack synchronization time $T_{\text{sy}}$ is shown in Table II.
TABLE II: Attack success probability (%) under different SIR and attack synchronization time $T_{\text{sy}}$
@@
Figure 6: Missing detection probability and false alarm probability of the attack detection scheme
Table II shows that the attack success probability decreases significantly when the attack synchronization time exceeds a certain range because it affects the detection of the SFD field and the demodulation of the PHR field at the receiver. According to the IEEE 802.15.4z protocol, there is a 128-bit gap between the front and the end of the STS segment, which can attenuate the impact of the attack signal on the SFD field detection and PHR field demodulation. Moreover, the peak value generated by the attack signal in the cross-correlation spectrum will be lower than the thresholds of the leading edge detection algorithm if the power of the STS segment of the attack signal is too low and the strongest path obtained by the leading edge detection algorithm may be changed if the power of the STS segment of the attack signal is too high, and attacks will fail in either of these circumstances.
To evaluate the performance of the proposed attack detection scheme in DS-TWR, we simulate under the Ghost Peak attack condition to obtain the missing detection probability and false alarm probability of the attack detection scheme with respect to SNR. As shown in Fig. 6, the attack detection scheme has a relatively low probability of missing detection and false alarm, and is robust to SNR. The reliability of the attack detection scheme provides a solid foundation for applying the random time-hopping DS-TWR.
The variation curve of attack success probability with SNR before and after using the classical DS-TWR and time-hopping DS-TWR under different SNR and SIR is shown in Fig. 7, where the random time-hopping delay $\Delta t$ is set to be from 15 to 20 microsecond.
@@
Fig. 7 indicates that the time-hopping DS-TWR can reduce the attack success probability to 0% in the simulation, reaching the order of one in ten thousand regardless of the attack success probability under normal circumstances. The proposed security strategy is robust to both SIR and SNR, verifying that the anti-attack mechanism performs well in resisting attacks.
@@ IV-B Experimental Validation
We validate the proposed security ranging strategy based on random time-hopping mechanism using commercial ranging devices embedded with Qorvo DW3110 chip in the Ghost Peak attack scenario (an indoor 10 m environment shown in Fig. 8). The results of the attack success probability are shown in Table. III with $\Delta t$ set in the range of -2 to 3600 microsecond in the Mode1 mode and the more vulnerable SDC mode of STS[16].
@@
According to listed results, the measurement of the chips can remain stable for tens of hours after adopting the proposed secure ranging strategy without redundant signaling overhead, which indicates that the proposed strategy performs satisfactorily in the practical environment.
@@ V Conclusion
In order to improve the security performance of existing UWB ranging, this paper proposes the time-hopping DS-TWR which is able to significantly reduce the attack success probability of various types of distance attacks represented by the Ghost Peak attack by increasing the randomness of the time of transmitting legitimate messages in the ranging process. Furthermore, combined with the attack detection scheme, the proposed secure ranging strategy is able to adapt to various UWB ranging application scenarios with very low complexity. However, this scheme mainly increases the randomness of the transmit time of the ranging message, and future work will incorporate more aspects (e.g. power and frame format) to optimize the security performance of the proposed strategy.

### v2

@@ III-C Performance Analysis
It is generally assumed that $t$ follows a uniform distribution $\text{U}(\text{T}_{\text{min}},\text{T}_{\text{max}})$, where $\text{T}_{\text{min}}$ and $\text{T}_{\text{max}}$ represent the minimum and maximum offsets of possible TOF in practice respectively. Taking into consideration (5) and (6), the attack success probability ${\mathbb{P}_{\text{s}}}^{{}^{\prime}}(t)$ at the time offset $t$ becomes:
${\mathbb{P}_{\text{s}}}^{{}^{\prime}}(t)$ $={\mathbb{P}_{\text{s}}}(t)~{}\mathbb{P}({T_{\text{SFD}}\leq t\leq T_{\text{payload}}})$ (7)
$=2\mathbb{P}\left({X>0.5\frac{\theta}{x_{t}}}+0.5\text{N}\right)\frac{\Delta t_{0}}{\Delta t_{1}},$
@@
$T_{\text{SFD}}\leq t-\Delta t\leq T_{\text{payload}}.$ (8)
Let the random variable $Y=t-\Delta t$. The attack success probability ${\mathbb{P}_{\text{s}}}^{{}^{\prime\prime}}(t)$ after adopting the random time-hopping mechanism becomes:
${\mathbb{P}_{\text{s}}}^{{}^{\prime\prime}}(t)$ $={\mathbb{P}_{\text{s}}}(t)~{}\mathbb{P}({T_{\text{SFD}}\leq t-\Delta t\leq T_{\text{payload}}})$ (9)
[…]
where ${f_{Y}}(y)$ is the probability density function (pdf) of the random variable $Y$, obtained by convolving two uniform distributions.
Let $\Delta t_{2}=T_{\text{max}}^{\text{hop}}-T_{\text{min}}^{\text{hop}}$, then ${f_{Y}}(y)$ is an isosceles trapezoid between $T_{\text{max}}^{\text{hop}}$ and $T_{\text{min}}^{\text{hop}}$, with a height of $\frac{1}{\Delta t_{2}}$ and an upper base of $\Delta t_{2}-\Delta t_{1}$. Since the range of delay $\Delta t_{2}$ for the proposed random time-hopping is much larger than the original possible TOF $\Delta t_{1}$ (i.e. $\Delta t_{2}\gg\Delta t_{1}$), ${f_{Y}}(y)$ can be approximated as a rectangle with bottom $\Delta t_{2}$ and height $\frac{1}{\Delta t_{2}}$. Therefore, the reduction factor $G$ of the attack success probability, representing the gain of this strategy, is:
@@ IV-A Numerical Simulation Evaluation
The signal parameter configurations for simulation are detailed in Table I. The processing flow of the receiver used in this paper includes shaping filtering, downsampling, cross-correlation, timestamp estimation and data demodulation employing a RAKE receiver architecture. The length of BTW is fixed at 400 samples, with a sampling rate of 2 GHz. The MPEP and PAPR thresholds are set to 0.5 and 2, respectively. Moreover, the distance between two legitimate devices is 10 m and an attack is considered successful when the measurement result is less than 5 m. A total of 20,000 DS-TWR simulations under the Ghost Peak attack are performed for each case.
@@
Payload encoding RS &Conv. Codes RS &Conv. Codes
@@
SIR is used to denote the ratio between the power of the STS segment of the legitimate signal and that of the attack signal, while SNR indicates the ratio between the power of the UWB signal and noise. The simulated attack success probability under different SIR and attack synchronization time $T_{\text{sy}}$ for a typical fixed SNR of -10 dB is shown in Table II.
TABLE II: Simulated Attack Success Probability
@@
Figure 5: Missing detection probability and false alarm probability of the attack detection scheme[14]
Table II shows that the attack success probability decreases significantly when the attack synchronization time exceeds a certain range because it affects the detection of the SFD field and the demodulation of the PHR field at the receiver. According to the IEEE 802.15.4z protocol, a 128-bit gap exists between the front and the end of the STS segment, which helps attenuate the impact of the attack signal on the SFD field detection and PHR field demodulation. Moreover, if the power of the STS segment of the attack signal is too low, the peak value in the cross-correlation spectrum will fall below the thresholds of the leading edge detection algorithm. Conversely, if the power is too high, the strongest path identified by the leading edge detection algorithm may be changed. Either of these two circumstances will cause attacks to fail.
To evaluate the performance of the proposed attack detection scheme in DS-TWR, simulations are conducted under the Ghost Peak attack condition to obtain the missing detection probability and false alarm probability of the attack detection scheme with respect to SNR. As shown in Fig. 5, the attack detection scheme has a relatively low probability of missing detection and false alarm, and is robust to variations in SNR. The reliability of the attack detection scheme provides a solid foundation for applying the random time-hopping DS-TWR.
The variation curve of attack success probability with SNR before and after using the classical DS-TWR and time-hopping DS-TWR under different SNR and SIR is shown in Fig. 6, where the random time-hopping delay $\Delta t$ is set between 15 and 20 µs.
@@
Fig. 6 indicates that the time-hopping DS-TWR can reduce the attack success probability to 0% in the simulation involving 20,000 attacks, reaching the order of one in ten thousand regardless of the attack success probability under normal circumstances. This result is consistent with the derived gain $G$. The proposed security strategy is robust to variations in both SIR and SNR, verifying that the anti-attack mechanism performs effectively in resisting attacks.
@@ IV-B Experimental Validation
We validate the proposed security ranging strategy based on random time-hopping mechanism using commercial ranging devices embedded with Qorvo DW3110 chip in a Ghost Peak attack scenario conducted in an indoor 10 m Line Of Sight (LOS) environment shown in Fig. 7. The attacker is deployed near the Initiator to attack the Response message. The results of the attack success probability are shown in Table III, with $\Delta t$ set in the range of -2 to 3600 µs in the Mode1 mode and the more vulnerable Super Deterministic Codes (SDC) mode of STS[17].
@@
According to the listed results, the measurement of the chips can remain stable for tens of hours after adopting the proposed secure ranging strategy without redundant signaling overhead, which indicates that the proposed strategy performs resiliently in the practical environment.
@@ V Conclusion
In order to improve the security performance of existing UWB ranging, this paper proposes the time-hopping DS-TWR which is able to significantly reduce the attack success probability of various types of distance attacks represented by the Ghost Peak attack by increasing the randomness of the time of legitimate message transmission during the ranging process. Furthermore, combined with the attack detection scheme, the proposed secure ranging strategy is resilient by adpating to various UWB ranging application scenarios with very low complexity. However, this scheme primarily increases the randomness of the transmit time of the ranging message, and future work will incorporate more aspects (e.g. frame format and large ranging networks) to further optimize the security performance of the proposed strategy. Considering the current applications of UWB in automotive digital keys, mobile robots, UAV collaborative positioning, and embodied navigation[18], further research on UWB secure ranging methods is expected in the presence of malicious interference or attacks.